OpenStudy (anonymous):

Given that f(x)=x+3 and g(x) = x^2-x, Find (f+g)(5), if it exists.

OpenStudy (rajee_sam):

(f + g ) just add the two functions. (f + g) = (x + 3 ) + ( x² - x) = x + 3 + x² - x = x² +3 now (f + g) (5) means vale of (f+g) at x = 5 (f+g) = x² + 3 (f + g) (5) = 5² + 3 = 28

OpenStudy (anonymous):

thank you